You have a circular garden with an area of 32 square feet. If you increase the radius by a scale factor of 5, what is the area of the new garden?

I think it may be:
32=r^2(pi)
so with scale factor of 5, it would be:
(5r)^2
so 25*32 or 800 square feet

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

You have a circular garden with an area of 32 square feet. If you increase the radius by a scale factor of 5, what is the area of the new garden?

I think it may be:
32=r^2(pi)
so with scale factor of 5, it would be:
(5r)^2
so 25*32 or 800 square feet
====================
800 sf is correct.
Area is a function of the square of the radius.
5^2 = 25
25*32 = 800

Answer by jim_thompson5910(35256)

You are correct. The answer is 800 square feet. The linear dimension increases by a scale factor of 5, so the area increases by a scale factor of 5^2 = 5*5 = 25. This idea works for any shape. It does not have to be a circle.